# A RankNet-Inspired Surrogate-Assisted Hybrid Metaheuristic for Expensive Coverage Optimization

### v1

@@ IV-A1 Initialization
To ensure a uniform exploration of the mixed-variable solution space, we use Sobol sampling [14] for discrete variables and Latin Hypercube Sampling [15] for continuous variables, respectively. A total of $2D$ samples are generated and evaluated using the real fitness function to initialize the data archive $\mathcal{A}$, where $D$ indicates the dimension of decision variables. The top-ranked $N$ individuals are then selected to form the initial population [16], where $N$ is the population size.
@@ IV-A2 Fitness Diversity-Driven Switching Strategy
We employ the fitness diversity metric proposed in [1] to guide the switching between global and local optimization. When current diversity falls below the threshold $\delta$, the alternative strategy is activated to discover diverse genotypes.
@@ IV-A3 RankNet-Based Pairwise Global Surrogate
For global optimization, we propose a novel global surrogate based on the RankNet framework [17], which employs pairwise training to rank individuals through parent-offspring comparison. Initially, the global surrogate is trained using all data from the archive $\mathcal{A}$. To reduce computational cost, the model is updated only when over $T$ new individuals are added to $\mathcal{A}$ and the global optimum improves. These new individuals are paired with old ones to generate data pairs for model updating; otherwise, the existing model is used for inference. Additionally, we design a data augmentation technique to enhance generalization for our deployment problem. For optimization, we adopt our proposed $CAGA_{mv}$ [5] to generate new individuals. The surrogate preselects $\tau$ individuals for real fitness evaluation, which are then added to $\mathcal{A}$.
@@ IV-A4 Surrogate-Assisted Local Estimation of Distribution
For local optimization, a distance-based RBFN is constructed as the local surrogate, while the weighted EDA [6] serves as the local optimizer. The RBFN is trained with the top-ranked $5D$ samples from the archive $\mathcal{A}$ [7]. After weighted EDA samples offspring, the best predicted and the most uncertain solutions are preselected for real fitness evaluation, and subsequently added to $\mathcal{A}$
@@ IV-A5 Termination
When the number of real FEs exceeds $MaxFEs$, the algorithm outputs the best solution obtained so far. Otherwise, RI-SHM goes back to Step 2 and iteratively repeats the procedure.
@@ IV-B RankNet-based Pairwise Global Surrogate
In the global-local surrogate-assisted optimization framework, the global surrogate and its optimizer should effectively explore the high-dimensional mixed-variable solution space to identify promising regions. Building on our previous work [5], we adopt $CAGA_{mv}$, a variant of GA, as the global optimizer due to its effective correlation-aware mechanism. The iteration of GA relies on the selection through parent-offspring comparison, which aligns well with pairwise-based surrogate models. It drives the population evolution by predicting the rankings between pairs of individuals [19]. Compared to regression-based surrogates, pairwise-based models are more robust as they focus on relative ordering rather than precise value prediction.
Similarly, in recommendation systems, researchers score and rank candidate items based on user preferences to generate personalized recommendation lists [20]. RankNet is a classic ranking framework that uses neural networks to model ranking functions, converting an ordinal regression problem into a simpler binary classification task [17]. For example, if item $A$ is ranked higher than $B$, and $B$ is ranked higher than $C$, then $A$ must be ranked higher than $C$ to maintain consistency [21]. To enforce this condition, RankNet applies the sigmoid function to calculate the preference probability, which is based on the difference between the network outputs of input pairs.
@@
As discussed in Section III, there is a strong correlation between discrete and continuous variables in the coverage optimization task. Therefore, we apply the multi-head attention mechanism [22] to effectively capture the intricate relationships between variables. For individual $i$, we use linear projections to obtain the embedding $e_{i}$ which includes the discrete part $e_{d}$ and two continuous parts $e_{c_{a}}$ and $e_{c_{b}}$. We then fuse the corresponding information with the Hadamard product. The output of this layer is subsequently delivered to the following MLPs to enhance the capacity of representing nonlinearity. All activation functions in the network are applied with ReLU, and batch normalization is implemented with PyTorch’s default parameters. Meanwhile, we do not employ dropout, following a similar approach to that in [8]. The pipeline of the global surrogate can be summarized as follows.
@@ IV-B1 Initiate the Surrogate
To initiate the global surrogate, we construct a training set $\mathcal{T}$ using the evaluated solutions from the archive $\mathcal{A}$. Suppose $\mathcal{A}=\{a_{1},a_{2},...,a_{M}\}$, where $M$ is the number of individuals in archive $\mathcal{A}$. Each training data can be represented as $\{a_{i},a_{j},y\},i\neq j$, where $label$ denotes the ranking between $a_{i}$ and $a_{j}$. Specifically, $y=0$ indicates that $a_{i}$ is superior, and vice versa. In this way, we construct the $\mathcal{T}$ with $M(M-1)$ individual pairs.
Since the generalization ability of deep networks can benefit from larger datasets [23], we introduce a novel data augmentation method to expand the datasets. According to the encoding format, the representation of individuals possesses the permutation invariance. For example, shuffling the discrete part while adjusting the continuous variables accordingly preserves the fitness ranking among individuals. Following this intuitive thought, the training dataset can be expanded by at most $P_{D}^{D}$ times, which effectively enhances the network performance without significantly increasing training time. Here, we choose to expand the training set 10 times for simplicity.
@@ IV-B2 Voting-Based Preselection Strategy
Taking the GA variant as the global optimizer, offspring are generated iteratively. Our goal is to select promising solutions from the offspring for real fitness evaluation, based on the trained global surrogate model. This serves as a preselection strategy within the model management of SAEAs.
@@ IV-B3 Surrogate Update Mechanism
As the optimization proceeds, the population evolves within the solution space, which deviates from its initial distribution. Updating the surrogate model at proper time improves accuracy while reducing unnecessary computation. Unlike the surrogates (i.e., RBFN, Kriging) used in most SAEAs, our model is updated only when the number of newly evaluated individuals added to $\mathcal{A}$ exceeds a threshold $T$ and the global optimum is updated.
We define the set $New=\{a_{new,1},a_{new,2},...,a_{new,T}\}$ to include the newly evaluated individuals. To mitigate catastrophic forgetting and enrich the training data, we sample $T$ old individuals from the archive $\mathcal{A}$ with a certain probability, forming the set $Old=\{a_{old,1},a_{old,2},...,a_{old,T}\}$ accordingly. These old and new individuals are paired as $[a_{old,i},a_{new,j}]$, which generates $T^{2}$ unseen pairs for model updating. Since the training data during updating is particularly sparse, we also apply the data augmentation technique to enlarge the training data. Due to the archive $\mathcal{A}$ has a capacity limit, we sample only from the most recent $T_{max}$ individuals to support online learning. During the update phase, the model is trained for a fixed number of $epoch=10$.
@@ IV-C Surrogate-Assisted Local Estimation of Distribution
The details of the algorithm are provided in Algorithm 1. Lines 1–4 compute the fitness weights for EDA and define the associated statistical model. In line 4, $I(a,b)$ is set to 1 if $a=b$, and 0 otherwise. Lines 6–13 describe the sampling procedure for both discrete and continuous variables, which are then combined to form $X$. [$\cdot$] indicates the operation of concatenation, and $N_{d}+N_{c}$ equals the total dimension $D$. In line 14, the trained RBFN is used to preselect $\tau$ individuals: the best individual is selected for exploitation, while the one with the greatest Euclidean distance from the current population, indicating the highest uncertainty, is chosen for exploration. In other words, here $\tau=2$.
@@ IV-D Fitness Diversity-Driven Switching Strategy
In the RI-SHM framework, global and local optimizations play distinct roles to complement each other. Global optimization focuses on exploring the mixed-variable solution space, while local optimization aims to refine promising solutions or escape from local optima when the global optimum stagnates. To push the solution towards better boundaries, this subsection introduces a switching strategy that effectively balances global and local optimization.
Population diversity is commonly used as an indicator for adaptively switching between global and local optimization. However, this metric may not be suitable for our deployment problem. Fig. 5 illustrates a top-down view showing two solutions with similar fitness values. However, they differ significantly in the solution space, especially in their discrete components and directional orientations. The experimental results also reveal that the population diversity metric for MVOPs, as proposed by [2], remains relatively high even during evolutionary stagnation. This highlights the importance of capturing fitness diversity in coverage optimization with multiple plateaus.
@@
Although fitness diversity may not accurately reflect the status of population, it provides an efficient way to decide the correct moment for switching between different strategies. When the fitness diversity is low, solutions are either clustered in a small region of the solution space, or distributed across plateaus with similar performance. In such cases, the activation of an alternative search can increase the chances to detect fresh genotypes [4]. In this paper, we set the threshold $\delta$ at $0.2$. If the fitness diversity falls below this value, we switch the optimizer; otherwise, we maintain the current one.
@@ IV-E Computational Complexity
During the iteration, the computational cost mainly arises from constructing the RBFN in the local optimization process and the online learning updates of the global surrogate model. Since the RBFN does not require an iterative training process,its time complexity is typically $O(K\cdot N_{sample})$, where $K$ stands for the number of basis functions and $N_{sample}$ is the number of samples. For simplicity, this component is ignored as $K<<N_{sample},N_{sample}<<M^{2}$. The training set for the global surrogate’s online learning is derived from real evaluated individuals, which are preselected within the overall framework under limited budget $MaxFEs$. Based on the update criterion, we assume that exactly $T$ new individuals are added to the archive $\mathcal{A}$ when the global optimum is updated, the training set contains $T^{2}$ data pairs. The model can be updated up to $MaxFEs/T$ times, with each individual trained for constant epochs. Thus, the computational complexity of this component is $O(T\cdot MaxFEs\cdot epoch)$.
@@ V Experimental results
In this section, we first outline the basic experimental settings. Next, we evaluate and validate the performance of RI-SHM across small to large scale deployment test instances. Finally, we conduct ablation studies to verify the effectiveness of RI-SHM’s core components.
@@ V-A1 Test Instances
Due to the inherent combinatorial property, the subset selection decision becomes increasingly complex as the number of candidate locations grows. Similarly, higher sampling precision and a greater number of targets further complicate the challenge of maximizing coverage.
@@
  • Notably, MixedEGO is excluded from large scale comparisons as it fails to solve these instances within 7 days due to high computational costs.
@@ V-A2 Algorithms in Comparison
    SHEDA [6]: An efficient algorithm belongs to EC-based optimization, which is frequently used as comparative algorithm in EMVOPs. It combines EDA with canonical Kriging to solve NAS.
@@
    DEDSO [50]: The latest algorithm [^4] https://github.com/ForrestXie9/DEDEO that presents a SAEA specifically for MVOPs, using $ACO_{mv}$ [52] as the base optimizer and RBFN as the primary surrogate. It exhibits strong performance in problems with dimensions up to 100.
@@ V-A3 Parameter Settings
In all algorithms compared, $2D$ samples are generated to initialize the data archive $\mathcal{A}$, from which the best $N$ individuals form the initial population. Here, $N$ is set to 100. The $MaxFEs$ is limited to 2000, which is only one-tenth of the original setting in [5] and aligns with the setting in [53].
@@ V-B1 Comparative Analysis of SAEAs
As shown in Table IV, RI-SHM consistently outperforms all comparative algorithms across the tested instances. Its advantages in stability and performance become more evident as problem scale grows. Fig. 8 illustrates the convergence curves of comparative algorithms across 15 instances.
In small and medium-scale scenarios, MixedEGO achieves results comparable to RI-SHM on certain instances, reflecting the strength of its kernel. However, RI-SHM exhibits a stronger ability to escape from local optima and continues to refine solutions during the later stages of evolution.
SHEDA demonstrates rapid convergence within 1,000 FEs in small and medium-scale problems, while fails to converge in large-scale scenarios, which indicates its strong exploratory potential. In contrast, DEDSO shows slower convergence in small and medium-scale problems and stagnates in large-scale scenarios. This indicates that DEDSO is more suitable for problems with fewer than 100 dimensions.
In large-scale scenarios, RI-SHM stands out by efficiently exploring the solution space within limited FEs. This is benefited from the pairwise global surrogate’s ability to quickly identify promising solutions. In comparison, algorithms like SHEALED, which perform well in small-scale problems and exhibit stagnation as problem complexity increases. These findings highlight RI-SHM’s scalability and effectiveness in addressing high-dimensional MVOPs.
@@ V-B2 Comparisons of RI-SHM and $CAGA_{mv}$
We compare RI-SHM with two versions of $CAGA_{mv}$ [5] in Table V. Under the same number of FEs, RI-SHM significantly outperforms $CAGA_{mv}$, highlighting the effectiveness of our SAEA framework. The smaller standard deviation indicates RI-SHM’s ability to consistently achieve high-quality solutions by balancing global and local optimization phases. Furthermore, Fig. 9 presents that even with increasing problem complexity, RI-SHM achieves comparable coverage performance to $CAGA_{mv}$-20K with slightly longer time. Here, the coverage rate is calculated as $Coverage=fitness/\sum_{q=1}^{|Q|}{\omega_{q}}$.

### v2

@@ IV-A1 Initialization
To ensure a uniform exploration of the mixed-variable solution space, we use Sobol sampling for discrete variables and Latin Hypercube Sampling for continuous ones, respectively. A total of $2D$ samples are generated and evaluated with a real fitness function to initialize the data archive $\mathcal{A}$, where $D$ indicates the dimension of decision variables. The top-ranked $N$ individuals are then selected to form the initial population [43], where $N$ is the population size.
@@ IV-A2 Fitness Diversity-Driven Switching Strategy
We employ the fitness diversity metric to guide the switching between global and local optimization. When current diversity falls below the threshold $\delta$, the alternative strategy is activated to discover diverse genotypes.
@@ IV-A3 RankNet-Based Pairwise Global Surrogate
For global optimization, we propose a novel global surrogate based on the RankNet framework [28], which employs pairwise training to rank individuals through parent-offspring comparison. Initially, the global surrogate is trained using all data from the archive $\mathcal{A}$. To reduce the computational cost, the model is updated only when over $T$ new individuals are added to $\mathcal{A}$ and the global optimum improves. These new individuals are paired with old ones to generate data pairs for model updating; otherwise, the existing model is used for inference. Additionally, we design a data augmentation technique to enhance generalization for our deployment problem. For optimization, we adopt our proposed $CAGA_{mv}$ [8] to generate new individuals. The surrogate preselects $\tau$ individuals for real fitness evaluation, which are then added to $\mathcal{A}$.
@@ IV-A4 Surrogate-Assisted Local Estimation of Distribution
For local optimization, a distance-based RBFN is constructed as the local surrogate, while the weighted EDA [18] serves as the local optimizer. The RBFN is trained with the top-ranked $5D$ samples from the archive $\mathcal{A}$ [11]. After weighted EDA samples offspring, the best predicted and most uncertain solutions are preselected for real fitness evaluation and added to $\mathcal{A}$
@@ IV-A5 Termination
When the number of real FEs exceeds $MaxFEs$, the algorithm outputs the best solution during the evolution. Otherwise, RI-SHM goes back to Step 2 and iteratively repeats the procedure.
@@ IV-B RankNet-based Pairwise Global Surrogate
In the global-local surrogate-assisted optimization framework, the global surrogate and its optimizer should effectively explore the high-dimensional mixed-variable solution space to identify promising regions. Building on our previous work [8], we adopt $CAGA_{mv}$, a variant of GA, as the global optimizer due to its effective correlation-aware mechanism. The iteration of GA relies on the selection through parent-offspring comparison, which aligns well with pairwise-based surrogate models. It drives population evolution by predicting the rankings between pairs of individuals. Compared to regression-based surrogates, pairwise-based models are more robust, focusing on relative ordering rather than precise value prediction.
Similarly, researchers score and rank candidate items based on user preferences in recommendation systems to generate personalized recommendation lists. RankNet is a classic ranking framework that uses neural networks to model ranking functions, converting an ordinal regression problem into a more straightforward binary classification task [28]. For example, if item $A$ is ranked higher than $B$, and $B$ is ranked higher than $C$, then $A$ must be ranked higher than $C$ to maintain consistency [44]. To enforce this condition, RankNet applies the sigmoid function to calculate the preference probability, which is based on the difference between the network outputs of input pairs.
@@
As discussed in Section III, there is a strong correlation between discrete and continuous variables in the coverage optimization task. Therefore, we apply the multi-head attention mechanism [45] to capture the intricate relationships between variables effectively. For individual $i$, we use linear projections to obtain the embedding $e_{i}$, which includes the discrete part $e_{d}$ and two continuous parts $e_{c_{a}}$ and $e_{c_{b}}$. We then fuse the corresponding information with the Hadamard product. The output of this layer is subsequently delivered to the following MLPs to enhance the capacity of representing nonlinearity. All activation functions in the network are applied with ReLU, and batch normalization is implemented with PyTorch’s default parameters. The pipeline of the global surrogate can be summarized as follows.
@@ IV-B1 Initiate the Surrogate
To initiate the global surrogate, we construct a training set $\mathcal{T}$ using the evaluated solutions from the archive $\mathcal{A}$. Suppose $\mathcal{A}=\{a_{1},a_{2},...,a_{M}\}$, where $M$ is the number of individuals in archive $\mathcal{A}$. Each training data can be represented as $\{a_{i},a_{j},y\},i\neq j$, where $label$ denotes the ranking between $a_{i}$ and $a_{j}$. Specifically, $y=0$ indicates that $a_{i}$ is superior and vice versa. In this way, we construct the $\mathcal{T}$ with $M(M-1)$ individual pairs.
Since the generalization ability of deep networks can benefit from larger datasets [46], we introduce a novel data augmentation method to expand the datasets. According to the encoding format, the representation of individuals possesses the permutation invariance. For example, shuffling the discrete part while adjusting the continuous variables accordingly preserves the fitness ranking among individuals. Following this intuitive thought, the training dataset can be expanded at most $P_{D}^{D}$ times, effectively enhancing the network performance without significantly increasing training time. Here, we choose to expand the training set 10 times for simplicity.
@@ IV-B2 Voting-Based Preselection Strategy
Taking the GA variant as the global optimizer, offspring are generated iteratively. We aim to select promising solutions from the offspring for real fitness evaluation based on the trained global surrogate model. This serves as a preselection strategy within the model management of SAEAs.
@@ IV-B3 Surrogate Update Mechanism
As the optimization proceeds, the population evolves within the solution space, which deviates from its initial distribution. Updating the surrogate model at the proper time improves accuracy while reducing unnecessary computation. Unlike the surrogates (i.e., RBFN, Kriging) used in most SAEAs, our model is updated only when the number of newly evaluated individuals added to $\mathcal{A}$ exceeds a threshold $T$ and the global optimum is updated.
We define the set $New=\{a_{new,1},a_{new,2},...,a_{new,T}\}$ to include the newly evaluated individuals. To mitigate catastrophic forgetting and enrich the training data, we sample $T$ old individuals from the archive $\mathcal{A}$ with a certain probability, forming the set $Old=\{a_{old,1},a_{old,2},...,a_{old,T}\}$ accordingly. These old and new individuals are paired as $[a_{old,i},a_{new,j}]$, which generates $T^{2}$ unseen pairs for model updating. Since the training data during updating is sparse, we also apply the data augmentation technique to enlarge the training data. Due to the archive $\mathcal{A}$ has a capacity limit, we sample only from the most recent $T_{max}$ individuals to support online learning. During the update phase, the model is trained for a fixed number of $epoch=10$.
@@ IV-C Surrogate-Assisted Local Estimation of Distribution
The algorithm details are provided in Algorithm 1. Lines 1–4 compute the fitness weights for EDA and define the associated statistical model. In line 4, $I(a,b)$ is set to 1 if $a=b$, and 0 otherwise. Lines 6–13 describe the sampling procedure for discrete and continuous variables, which are combined to form $X$. [$\cdot$] indicates the operation of concatenation, and $N_{d}+N_{c}$ equals the total dimension $D$. In line 14, the trained RBFN is used to preselect $\tau$ individuals: the best individual is selected for exploitation, while the one with the greatest Euclidean distance from the current population, indicating the highest uncertainty, is chosen for exploration. In other words, here $\tau=2$.
@@ IV-D Fitness Diversity-Driven Switching Strategy
In the RI-SHM framework, global and local optimizations play distinct roles that complement each other. Global optimization focuses on exploring the mixed-variable solution space, while local optimization aims to refine promising solutions or escape from local optima when the global optimum stagnates. To push the solution towards better boundaries, this subsection introduces a switching strategy that effectively balances global and local optimization.
Population diversity is commonly used as an indicator for adaptively switching between global and local optimization. However, this metric may not be suitable for our deployment problem. Fig. 5 illustrates a top-down view showing two solutions with similar fitness values. However, their discrete components and directional orientations differ significantly in the solution space. The experimental results also reveal that the population diversity metric for MVOPs, as proposed by [47], remains relatively high even during evolutionary stagnation. This highlights the importance of capturing fitness diversity in coverage optimization with multiple plateaus.
@@
Although fitness diversity may not accurately reflect the status of the population, it provides an efficient way to decide the correct moment for switching between different strategies. When the fitness diversity is low, solutions are clustered in a small region of the solution space or distributed across plateaus with similar performance. In such cases, activating an alternative search can increase the chances of detecting fresh genotypes [48]. In this paper, we set the threshold $\delta$ at $0.2$. If the fitness diversity falls below this value, we switch the optimizer; otherwise, we maintain the current one.
@@ IV-E Computational Complexity
During the iteration, the computational cost mainly arises from constructing the RBFN in the local optimization process and the online learning updates of the global surrogate model. Since the RBFN does not require an iterative training process, its time complexity is typically $O(K\cdot N_{sample})$, where $K$ stands for the number of basis functions and $N_{sample}$ is the number of samples. For simplicity, this component is ignored as $K<<N_{sample},N_{sample}<<M^{2}$. The training set for the global surrogate’s online learning is derived from real evaluated individuals, which are preselected within the overall framework under a limited budget of $MaxFEs$. Based on the update criterion, we assume that exactly $T$ new individuals are added to the archive $\mathcal{A}$ when the global optimum is updated, and the training set contains $T^{2}$ data pairs. The model can be updated up to $MaxFEs/T$ times, with each individual trained for constant epochs. Thus, the computational complexity of this component is $O(T\cdot MaxFEs\cdot epoch)$.
@@ V Experimental results
In this section, we outline the basic experimental settings. Next, we evaluate and validate the performance of RI-SHM across small to large-scale deployment test instances. Finally, we conduct ablation studies to verify the effectiveness of RI-SHM’s core components.
@@ V-A1 Test Instances
Due to the inherent combinatorial property, the subset selection decision becomes increasingly complex as the number of candidate locations grows. Similarly, higher sampling precision and a more significant number of targets further complicate the challenge of maximizing coverage.
@@
  • Notably, MixedEGO is excluded from large-scale comparisons as it fails to solve these instances within 7 days due to high computational costs.
@@ V-A2 Algorithms in Comparison
    SHEDA [18]: An efficient algorithm belongs to EC-based optimization, frequently used as a comparative algorithm in EMVOPs. It combines EDA with canonical Kriging to solve NAS.
@@
    DEDSO [22]: The latest algorithm [^4] https://github.com/ForrestXie9/DEDEO that presents an SAEA specifically for MVOPs, using $ACO_{mv}$ [16] as the base optimizer and RBFN as the primary surrogate. It exhibits strong performance in problems with dimensions up to 100.
@@ V-A3 Parameter Settings
In all algorithms compared, $2D$ samples are generated to initialize the data archive $\mathcal{A}$, from which the best $N$ individuals form the initial population. Here, $N$ is set to 100. The $MaxFEs$ is limited to 2000, only one-tenth of the original setting in [8] and aligns with the setting in [38].
@@ V-B1 Comparative Analysis of SAEAs
As shown in Table IV, RI-SHM consistently outperforms all comparative algorithms across the tested instances. Its advantages in stability and performance become more evident as the problem scale grows. Fig. 8 illustrates the convergence curves of comparative algorithms across 15 instances.
In small and medium-scale scenarios, MixedEGO achieves results comparable to RI-SHM in certain instances, reflecting the strength of its kernel. However, RI-SHM exhibits a stronger ability to escape from local optima and continues to refine solutions during the later stages of evolution.
SHEDA demonstrates rapid convergence within 1,000 FEs in small and medium-scale problems while failing to converge in large-scale scenarios, which indicates its strong exploratory potential. In contrast, DEDSO shows slower convergence in small and medium-scale problems and stagnates in large-scale scenarios. This indicates that DEDSO is more suitable for problems with fewer than 100 dimensions.
In large-scale scenarios, RI-SHM stands out by efficiently exploring the solution space within limited FEs. This is benefited from the pairwise global surrogate’s ability to identify promising solutions quickly. In comparison, algorithms like SHEALED perform well in small-scale problems and exhibit stagnation as problem complexity increases. These findings highlight RI-SHM’s scalability and effectiveness in addressing high-dimensional MVOPs.
@@ V-B2 Comparisons of RI-SHM and $CAGA_{mv}$
We compare RI-SHM with two versions of $CAGA_{mv}$ [8] in Table V. Under the same number of FEs, RI-SHM significantly outperforms $CAGA_{mv}$, highlighting the effectiveness of our SAEA framework. The smaller standard deviation indicates RI-SHM’s ability to consistently achieve high-quality solutions by balancing global and local optimization phases. Furthermore, Fig. 9 presents that even with increasing problem complexity, RI-SHM achieves comparable coverage performance to $CAGA_{mv}$-20K with a slightly longer time. Here, the coverage rate is calculated as $Coverage=fitness/\sum_{q=1}^{|Q|}{\omega_{q}}$.